In Youngs double experiment,the fringe width is β.If the entire arrangement is now placed inside a liquid of refractive index μ,the fringe width will become


a.μβ


b.β/μ


c.β/(μ+1)


d.β/(μ-1)

Grade:12th Pass

1 Answers

Arun
3 years ago
If the screen distance id D, distance between the slits is d and the wave length is λ
 
then the fringe width is given by
 
b = λD/d
 
When the arrangement is immerse in a liquid of refractive index u then the new wavelength is 1/u times the original wavelength
 
If the new wavelength is λ' then
 
λ' = λ/n
 
Therefore, the new fringe width will be
 
b' = λ'D/d
 
 
 
or, b' = b/u
 
Therefore the fringe width will decrease by a factor of u.
 
 
 
Thanks and regards